OpenStudy (anonymous):

Solve. Express your answer in simplest form. 2/5*5 A.7/25 B.1/2 C.1 2/5 D.2

OpenStudy (anonymous):

D

OpenStudy (anonymous):

When you convert 2/5 to a decimal you get .4 and 4*5=20 so you move over the decimal point because you are multiplying a decimal and get 2.0

Nnesha (nnesha):

you don't even need a calculator for this number hint \[\frac{y}{x} *x \rightarrow \frac{y}{\cancel{x}} * \cancel{x} \rightarrow y \]